OpenStudy (anonymous):

estimate the value of √35

OpenStudy (shane_b):

It's almost the same as \[\sqrt{36}\]which = 6...so a good estimate would be 6.

OpenStudy (anonymous):

about 5.9 because sqare of 6 is 36 and square of 5 is 25 so its much closer to 36

OpenStudy (shane_b):

Right...I should have wrote that it should be just under 6.

OpenStudy (lgbasallote):

is this linear approximation?

OpenStudy (anonymous):

thanks for your help guys, how can I proof it? I'm sorry I am week in math and I just want to know how you found it? Please help...

OpenStudy (anonymous):

no that is Root and Powers.

OpenStudy (lgbasallote):

oh darn i thought it was calculus

OpenStudy (anonymous):

nice profile picture though :)

OpenStudy (lgbasallote):

1^2 = 1 2^2 = 4 3^2 = 9 4^2 = 16 5^2 = 25 6^2 = 36 35 is in between the square of 5 and 6 so it's around 5 point something then it's VERY close to 36 so we can assume maybe 5.8 or 5.9

OpenStudy (anonymous):

Thank You :)

OpenStudy (lgbasallote):

very much welcome

OpenStudy (anonymous):

can you help me more on the assignment?